# Connecting the Dots: A Chain-of-Collaboration Prompting Framework for LLM Agents

### v1

@@ C.1 Efficiency and Cost Analysis
PMC [68] 241.56 11637.94 6644.22 0.0359
MedAgents [50] 196.53 3901.03 3483.24 0.0170
Debate(short) [20] 187.16 3938.35 4809.27 0.0223
Debate(long) [20] 162.57 4357.87 5278.53 0.0246
CoA [72] 227.21 11428.06 5446.62 0.0309
@@
PMC [68] 198.49 3796.07 2613.73 0.0135
MedAgents [50] 169.53 3504.85 3504.85 0.0168
Debate(short) [20] 548.66 3372.56 3772.79 0.0178
Debate(long) [20] 574.95 3810.87 4268.74 0.0201
CoA [72] 253.18 5493.94 3450.45 0.0182
@@
PMC [68] 256.84 11319.74 6452.89 0.0349
MedAgents [50] 181.61 3816.26 3445.34 0.0169
Debate(short) [20] 318.51 4414.23 5660.94 0.0261
Debate(long) [20] 293.42 5438.52 6911.85 0.0321
CoA [72] 195.49 11446.05 5353.24 0.0306
@@
C.2 More Experiments on Single Agent
Table 9 and Table 10 demonstrate that integrating Cochain markedly enhances single-agent performance across diverse backbone models during two distinct stages of a business workflow. This enhancement is consistently reflected across all evaluated metrics, irrespective of whether the LLMs are in their raw, pre-trained state or have undergone fine-tuning. Notably, a synergistic effect between fine-tuning and the application of Cochain typically yields optimal performance, underscoring Cochain’s robust capability to leverage specialized model knowledge. The broad-spectrum improvements observed across various model metrics highlight the effectiveness and versatility of Cochain in mitigating prior issues of undercollaboration.
@@ C.2 More Experiments on Single Agent
C.3 More Experiments on Specialized Agent Collaboration
Table 4 presents the results from a distinct experimental configuration designed to further investigate the coordination capabilities of Cochain. Whereas Table 9 assessed a multi-domain model (a single Qwen2-7B model fine-tuned on all domains) and specialized models (different models each fine-tuned for specific domains), Table 11 exclusively focuses on a setup utilizing multiple agents of the same backbone, each independently fine-tuned for a specific domain. The findings indicate that Cochain effectively facilitates collaboration among these identically architected yet uniquely specialized agents. Performance enhancements are observed both within their respective specialized domains and in the quality of the comprehensive, integrated solutions. This underscores Cochain’s distinct advantage in synergizing specialized knowledge from a team of uniquely trained agents that share a common architectural foundation.
Table 11: Results for one model at particular stages. This means that we use Qwen2-7B for four fine-tunings, and the four fine-tuned models are experimented with at their respective stages. We highlight the fine-tuned results, Cochain application, and improvements()($\uparrow$)( ↑ ).
@@ C.3 More Experiments on Specialized Agent Collaboration
C.4 More Results from Ablation Studies
On the automotive dataset, we report additional ablation study results focusing on the impact of different backbone models. As indicated in Table 12, “w/o Prompts Tree” markedly impaired performance. Specifically, when applied to the Llama-8B backbone, this ablation led to a 1.97% reduction in the BS-F score, with notable degradations also observed across other evaluation metrics. These findings are consistent with the experimental results presented in Table 5, further underscoring the critical role of the Prompts Tree component in our framework.
Table 12: More results of ablation study. $\mathcal{KG}$ represents the knowledge graph. We highlight the most and second-most efficient modules.
@@ C.4 More Results from Ablation Studies
C.5 Integration of Cochain with Baselines
We further investigate the integration of Cochain with existing baseline methods, specifically PMC and MedAgents, through experiments conducted on a pharmaceutical dataset using Claude-3.5-Haiku. As reported in Table 13, the combination of Cochain with PMC results in improvements of 2.62% and 7.55% in the BS-F and BLEU-4 scores, respectively. This outcome further substantiates the efficacy of Cochain in mitigating issues associated with over-collaboration.
@@ C.5 Integration of Cochain with Baselines
PMC [68] 66.51 66.06 66.97 5.17 12.33 20.31 5.48 13.56
PMC [68] + Cochain 69.13 68.49 69.80 12.72 18.48 27.01 6.93 15.82
MedAgents [50] 66.23 66.02 66.48 7.17 14.02 22.59 5.43 14.32
MedAgents [50] + Cochain 67.45 66.91 68.00 11.65 17.80 27.06 6.16 15.74
Appendix D Broader Impacts
Cochain enhances agent efficiency and effectiveness in complex business workflows, potentially increasing productivity and enabling agent systems to tackle more sophisticated multi-step tasks across various industries. By addressing both under-collaboration and over-collaboration, Chain-of-Collaboration contributes to more robust and reliable agent systems. The interpretability offered by our framework, for instance through the causal chain mechanism, can also aid in scrutinizing agent behavior, potentially reducing misuse. Cochain’s design focuses on improving collaboration quality rather than solely on automation. Nevertheless, the broader societal implications of increased automation in workflows, where more tasks are executed by agent systems, warrant ongoing consideration. For the responsible dissemination of our data, which is created using Self-Instruct and consists of generalized, non-sensitive information that does not contain Personally Identifiable Information or private commercial data, we will provide comprehensive documentation, a clear academic research license stipulating ethical use, and a designated point of contact for inquiries.

### v2

@@ D.1 More Efficiency and Cost Analysis
PMC Zhang et al. [2025a] 241.56 11637.94 6644.22 0.0359
MedAgents Tang et al. [2024] 196.53 3901.03 3483.24 0.0170
Debate(short) Du et al. [2024] 187.16 3938.35 4809.27 0.0223
Debate(long) Du et al. [2024] 162.57 4357.87 5278.53 0.0246
CoA Zhang et al. [2024b] 227.21 11428.06 5446.62 0.0309
@@
PMC Zhang et al. [2025a] 198.49 3796.07 2613.73 0.0135
MedAgents Tang et al. [2024] 169.53 3504.85 3504.85 0.0168
Debate(short) Du et al. [2024] 548.66 3372.56 3772.79 0.0178
Debate(long) Du et al. [2024] 574.95 3810.87 4268.74 0.0201
CoA Zhang et al. [2024b] 253.18 5493.94 3450.45 0.0182
@@
PMC Zhang et al. [2025a] 256.84 11319.74 6452.89 0.0349
MedAgents Tang et al. [2024] 181.61 3816.26 3445.34 0.0169
Debate(short) Du et al. [2024] 318.51 4414.23 5660.94 0.0261
Debate(long) Du et al. [2024] 293.42 5438.52 6911.85 0.0321
CoA Zhang et al. [2024b] 195.49 11446.05 5353.24 0.0306
@@
Table 8: Latency Breakdown Analysis of Cochain’s Inference Components
Inference Component Average Time (s) Percentage of Total Time (%)
Knowledge Graph Retrieval 0.6200 s 1.57 %
Prompts Tree Retrieval 0.0022 s <0.01 %
LLM Model Inference 38.7700 s 98.42 %
D.2 More Experiments on Single Agent
Table 9 and Table 10 demonstrate that integrating Cochain markedly enhances single-agent performance across diverse backbone models during two distinct stages of a business workflow. This enhancement is consistently reflected across all evaluated metrics, irrespective of whether the LLMs are in their raw, pre-trained state or have undergone fine-tuning. Notably, a synergistic effect between fine-tuning and the application of Cochain typically yields optimal performance, underscoring Cochain’s robust capability to leverage specialized model knowledge. The broad-spectrum improvements observed across various model metrics highlight the effectiveness and versatility of Cochain in mitigating prior issues of insufficient collaboration.
@@ D.2 More Experiments on Single Agent
Table 11: Quantifying insufficient collaboration.
Backbone Method BS-F GLEU ROUGE-L
Qwen2-7B Single Agent 71.72 21.64 24.07
+ Manual Prompts 72.54 ($\uparrow$) 21.83 ($\uparrow$) 24.34 ($\uparrow$)
+ Cochain 75.39 ($\Uparrow$) 28.17 ($\Uparrow$) 28.87 ($\Uparrow$)
DeepSeek-R1-7B Single Agent 71.35 19.55 22.71
+ Manual Prompts 71.83 ($\uparrow$) 19.90 ($\uparrow$) 23.38 ($\uparrow$)
+ Cochain 73.19 ($\Uparrow$) 21.71 ($\Uparrow$) 24.01 ($\Uparrow$)
As shown in Table 11, to better quantify the impact of insufficient collaboration, we compare the baseline single agent to one provided with minimal communication, simulated by manually injecting five prompts containing cross-domain constraints. The results of this table clearly show: First, insufficient collaboration is a real problem, as the agent with injected communication showed a slight performance improvement over the original agent. This directly proves that adding cross-domain communication improves outcomes, thereby quantifying the performance loss from insufficient collaboration. Second, Cochain’s intelligent collaboration is far superior to simple communication, as the performance improved substantially after applying Cochain. This demonstrates that Cochain’s value is not just about adding information. Rather, it excels at intelligently filtering, organizing, and structuring collaborative knowledge. Its effectiveness far surpasses that of unstructured, manually added information and helps to avoid the risk of excessive collaboration.
D.3 More Experiments on Specialized Agent Collaboration
Figure 6 (Main Paper) presents the results from a distinct experimental configuration designed to further investigate the coordination capabilities of Cochain. Whereas Table 9 assessed a multi-domain model (a single Qwen2-7B model fine-tuned on all domains) and specialized models (different models each fine-tuned for specific domains), Table 12 exclusively focuses on a setup utilizing multiple agents of the same backbone, each independently fine-tuned for a specific domain. The findings indicate that Cochain effectively facilitates collaboration among these identically architected yet uniquely specialized agents. Performance enhancements are observed both within their respective specialized domains and in the quality of the comprehensive, integrated solutions. This underscores Cochain’s distinct advantage in synergizing specialized knowledge from a team of uniquely trained agents that share a common architectural foundation.
Table 12: Results for one model at particular stages. This means that we use Qwen2-7B for four fine-tunings, and the four fine-tuned models are experimented with at their respective stages. We highlight the fine-tuned results, Cochain application, and improvements()($\uparrow$).
@@ D.3 More Experiments on Specialized Agent Collaboration
D.4 More Results from Ablation Studies
On the automotive dataset, we report additional ablation study results focusing on the impact of different backbone models. As indicated in Table 13, “w/o Prompts Tree” markedly impaired performance. Specifically, when applied to the Llama-8B backbone, this ablation led to a 1.97% reduction in the BS-F score, with notable degradations also observed across other evaluation metrics. These findings are consistent with the experimental results presented in Table 5 (Main Paper), further underscoring the critical role of the Prompts Tree component in our framework.
Table 13: More results of the ablation study. $\mathcal{CKG}$ represents the collaborative knowledge graph. We highlight the most and second-most efficient modules.
@@ D.4 More Results from Ablation Studies
D.5 Integration of Cochain with Baselines
We further investigate the integration of Cochain with existing baseline methods, specifically PMC and MedAgents, through experiments conducted on a pharmaceutical dataset using Claude-3.5-Haiku. As reported in Table 14, the combination of Cochain with PMC results in improvements of 2.62% and 7.55% in the BS-F and BLEU-4 scores, respectively. This outcome further substantiates the efficacy of Cochain in mitigating issues associated with excessive collaboration.
@@ D.5 Integration of Cochain with Baselines
PMC Zhang et al. [2025a] 66.51 66.06 66.97 5.17 12.33 20.31 5.48 13.56
PMC Zhang et al. [2025a] + Cochain 69.13 68.49 69.80 12.72 18.48 27.01 6.93 15.82
MedAgents Tang et al. [2024] 66.23 66.02 66.48 7.17 14.02 22.59 5.43 14.32
MedAgents Tang et al. [2024] + Cochain 67.45 66.91 68.00 11.65 17.80 27.06 6.16 15.74
D.6 Evaluation of Collaborative Knowledge Graph
To evaluate the reliability of our final knowledge graph, we randomly sampled 2,000 triplets from it, consisting of 1,000 from explicit knowledge extraction and 1,000 from our counterfactual method. We used three core metrics for evaluation: Factual Correctness (Is the knowledge factually correct in the real world?),Task Relevance (How helpful is the knowledge for solving business workflow problems?), and Knowledge Depth (Is the knowledge common sense, or does it require professional insight?) We invited two domain experts for a blind review. We also used Gemini-2.5-Pro as a judge for an objective evaluation. The average scores (out of 5) are presented in the Table 15.
Table 15: Evaluation of Collaborative Knowledge Graphs by Human Experts and LLMs.
Evaluation Metric Human Expert Score (Avg. / 5.0) Gemini-2.5-Pro Score (Avg. / 5.0)
Factual Correctness 4.99 5.00
Task Relevance 4.76 4.81
Knowledge Depth 4.31 4.54
D.7 Excessive collaboration is More Serious than Insufficient collaboration
Business workflow tasks are particularly susceptible to the “excessive collaboration” phenomenon. When multiple agents collaborate, this manifests as responses deviating from core issues and reducing answer quality. Comparative analysis of Table 1 (Main Paper), Figure 6 (Main Paper), and additional single-agent experimental results (Supplementary Material D.2) demonstrates that, despite consuming significantly more computational resources, excessive collaboration performs substantially worse than insufficient collaboration. excessive collaboration’s challenging nature is its covertness—systems maintain high activity and apparent collaboration, making the problem hard to detect and rectify promptly. While task decomposition, such as approaches like PMC, and summarization, such as strategies like CoA, are considered effective methods for focusing on core issues Wu et al. [2024]; Chen et al. [2024], Cochain’s collaboration mechanism exhibits superior performance in controlling excessive collaboration, enabling a more precise focus on critical task requirements.
D.8 Robustness to Incomplete Retrieval
To validate the system’s robustness against potential keyword retrieval failures, we conducted a stress test. We designed 50 structured queries in the format of “How to use [Core Concept A] to solve problems related to [Target Concept B]”. In the test condition, we deliberately masked all keywords corresponding to [Target Concept B] to simulate a first-stage retrieval failure. For instance, in the query “How can ergonomics be used to reduce traffic accident injuries?”, the keywords “traffic” and “accident” were blocked.
While end-to-end semantic retrieval could theoretically solve keyword-mismatch issues, it becomes computationally prohibitive for real-time interaction on massive knowledge graphs. We therefore adopt the Retrieval-Ranking two-stage paradigm, a proven industry standard Covington et al. [2016]. This approach combines the advantages of both methods: keyword-based retrieval in the first stage ensures precise matching of technical terms and proper nouns, avoiding omissions from semantic over-generalization. In the second stage, semantic ranking captures the query’s overall intent, ensuring system generalization and robustness.
Table 16: Performance comparison under simulated retrieval failure. Masking the target concept’s keywords results in a negligible performance drop, demonstrating the system’s robustness.
Test Condition BS-F GLEU ROUGE-L
Full Query 74.16 22.08 25.86
Target Concept Masked 73.87 21.42 24.57
Performance Drop (%) -0.29% -0.66% -1.29%
The results in Table 16 confirm the effectiveness of this design. Even when forcing a first-stage failure, the BS-F score experienced a minimal drop of only 0.29%, from 74.16 to 73.87, with similarly small decreases in GLEU (-0.66%) and ROUGE-L (-1.29%) scores. This resilience is attributed to two factors. First, as justified above, the semantic ranking stage corrects for an imperfect initial candidate set. More critically, the subsequent graph traversal and causal chain stages provide a powerful secondary path to completeness. Even if “Traffic Accident” is missed during retrieval, it can be discovered through pre-existing knowledge paths represented as triples (e.g., (Injuries, is_a_result_of, Traffic Accident)), ensuring the final context remains robust and comprehensive.